Please use the following information to answer the remaining problems: Able Corporation has a project with the following cash flows and an 9.6% cost of money: Numbers in parentheses are outflows. Both Year 0 and Year 3 cash flows are outflows. Year 0 1 2 3 4 5 6 Cash flow $(351,000) $ 95,000 $186,000 $(300,000) $ 280,000 $260,000 $268,000 23. Please calculate the net present value $175,730.78 24. Please calculate the profitability indexes (two decimals please) 1.50 25. Please calculate the modified profitability index using the terminal value approach in the textbook (two decimals please) 1.20 26. Please calculate the internal rate of return (two decimals please) 21.13% 27. Please calculate the modified internal rate of return (two decimals please and per the book)________________________ 28. Please calculate the payback period (two decimals please)________________________ 29. Please calculate the present value payback period (two decimals please)______________________
Answer #1

Since the question involves massive calculations, I am using excel.

H с 0 E Cumulative Discount Cumulative Cash flows DCF Cash flows factor DCF Cumulative CDF D= CxDF Cumulative D (351,000) (35

You already have got the correct answers for the earlier questions. Hence, I am not addressing them.

27. Please calculate the modified internal rate of return (two decimals please and per the book) = 14.55% (Please see the table above, adjacent cell contains the formula in excel used to get the answer)

28. Please calculate the payback period (two decimals please): Please see the cumulative cash flows. Please see the cells highlighted in yellow color. Hence, payback period = 4 + 90,000 / 260,000 = 4.35 years

29. Please calculate the present value payback period (two decimals please): Please see the cumulative DCFs. Please see the cells highlighted in green color. Hence, present value payback period = 4 + 143,298.39 / 164,407.10 = 4.87 years

= 4.35 years